OpenStudy (anonymous):

6(x-9)= 9 (x+4 ) solving proportions

OpenStudy (anonymous):

6x-54=9x+36 -3x=90 x=-30

OpenStudy (anonymous):

6x-54=9x+36 -3x=90 x- you always beat me

OpenStudy (anonymous):

thanks so much